OpenStudy (anonymous):

Solve the equation. k – 5 = –2

myininaya (myininaya):

k-5=-2 k-5+5=-2+5 k=3

OpenStudy (anonymous):

Add 5 to each side -2 + 5 = 3 k = 3

OpenStudy (anonymous):

k=3